The second hand of a clock is 8.4 centimeters long. Determine the linear speed of the tip of the second hand as it moves through 1 revolution around the clock face (using SI units).
Answer #1

Liner speed = distance/ time

Diameter = 2r = 2 (1.50) = 3.00cm or 0.03m

Circumference = pi x Diameter

Circumference = 3.14 x 0.03m

Circumference = 0.0942m
for one revolution is 60sconds
Linear speed = Distance / time

Linear speed = 0.0942/60 =1.57x10-3s